Oxygen Optimization is considering buying a new purification system. The new system would be purchased today for 19,600 dollars. It would be depreciated straight-line to 1,000 dollars over 2 years. In 2 years, the system would be sold and the after-tax cash flow from capital spending in year 2 would be 2,300 dollars. The system is expected to reduce costs by 7,400 dollars in year 1 and by 15,600 dollars in year 2. If the tax rate is 50 percent and the cost of capital is 10.1 percent, what is the net present value of the new purification system project?
Answer #1

rate positively ..

Annual depreciation = (19600-1000)/2 9300
year 0 1 2
A Initial investment -19600
operating cash flow
i Saving in cost        7,400        15,600
ii depreciation        9,300           9,300
iii=i-ii Taxable income       (1,900)           6,300
iv=iii*50% Tax@ 50%          (950)           3,150
v=iii-iv Profit after tax          (950)           3,150
B=v+ii operating cash flow after tax        8,350        12,450
C Post tax salvage value =           2,300
D=A+B+C Net cash flow after tax        (19,600)        8,350        14,750
E= PIVF @ 10.1% 1 0.908265 0.8249457
F=D*E Present value = (19,600.00) 7,584.01 12,167.95     151.96
Therefore NPV =          151.96